# FSGe: A fast and strongly-coupled 3D fluid-solid-growth interaction method

### v1

@@ 5.6 Pulsatility and directionality of blood flow
Our FSGe model considered steady-state flow conditions as an input for the intramural and WSS G&R stimuli, $\Delta\sigma_{I}$ and $\Delta\tau_{w}$, respectively. This is equivalent to some previous FSG methods in Table 1, whereas others used time-averaged G&R stimuli from a pulsatile flow simulation. Pulsatility is an important determinant of the blood vessel wall geometry, structure, and properties [72, 73]. For example, Eberth et al. [74] found that increased wall thickness in hypertension correlated more strongly with pulse than with mean or systolic pressure. Extending our FSGe approach to incorporate pulsatility would require solving an fluid-structure interaction (FSI) problem with an elastic wall in each coupling iteration. We could then extract the instantaneous elastic properties of the CMMe with a fixed G&R state from the evolved equilibrial state, as shown by Latorre and Humphrey [5]. In the aneurysmal region, we currently introduce dysfunctional mechanosensing at the endothelium by regionally setting $K_{\tau\sigma,o}\to 0$ in the top part of the aneurysm [56] as a model input. In the future, mechanosensing could be a function of pulsatility and thus as part of the model prediction.
@@
Another crucial input would be the variability of WSS over a cardiac cycle. Figure 14 shows a recirculation zone with backflow against the predominant flow direction at the top of the aneurysm. This flow separation only occurs in the FSGe model with $K_{\tau\sigma,o}=0.0$ (top left in Figure 12) due to its nearly two-fold increase in radius. Note, however, that the maximum backflow velocity is only 1.6% of the maximum inlet velocity. This effectively reverses the direction of the local WSS vector $\bm{\tau}_{w}$ acting on the vessel wall. However, in solid G&R models, $\tau_{w}$ can only be estimated as a scalar quantity to inform the WSS stimulus, typically derived from Poiseuille flow (1). It is well established that not only the magnitude but also the pulsatility and direction of WSS are constantly sensed by the endothelium [75, 76]. For example, the directionality of WSS plays an important role in the development of atherosclerosis [77]. Mohamied et al. [78] proposed a new metric,
@@
to quantify the change of direction of WSS over the course of a cardiac cycle $T$, with vessel normal vector $\bm{n}$ and mean WSS vector $\bm{\tau}_{\text{mean}}$. This metric was shown to correlate better with atherosclerotic lesion prevalence than other metrics like time-averaged or oscillatory WSS index. Our FSGe approach allows locally quantifying directionality measures like transWSS. In the future, incorporating local and directional WSS will enable the testing of hypotheses in FSGe-mediated aneurysm growth and atherosclerotic lesion formation using FSGe. Since current G&R models were designed with the static and scalar nature of reduced fluid dynamics in mind, we expect future biological insights from FSGe to arise from the advancements outlined in this work.

### v2

@@ 5.6 Pulsatility and directionality of blood flow
Our FSGe model considered steady-state flow conditions as an input for the intramural and WSS G&R stimuli, $\Delta\sigma_{I}$ and $\Delta\tau_{w}$, respectively. This is equivalent to some previous FSG methods in Table 1, whereas others used time-averaged G&R stimuli from a pulsatile flow simulation. Pulsatility is an important determinant of the blood vessel wall geometry, structure, and properties [79, 80]. For example, Eberth et al. [81] found that increased wall thickness in hypertension correlated more strongly with pulse than with systolic and especially mean blood pressure. Extending our FSGe approach to incorporate pulsatility would require solving an fluid-structure interaction (FSI) problem with an elastic wall in each coupling iteration. We could then extract the instantaneous elastic properties of the CMMe with a fixed G&R state from the evolved equilibrial state, as shown by Latorre and Humphrey [5]. In the aneurysmal region, we currently introduce dysfunctional mechanosensing at the endothelium by regionally setting $K_{\tau\sigma,o}\to 0$ in the top part of the aneurysm [63] as a model input. In the future, mechanosensing could be a function of pulsatility and thus as part of the model prediction.
@@
Another crucial input would be the variability of WSS over a cardiac cycle. Figure 14 shows a recirculation zone with backflow against the predominant flow direction at the top of the aneurysm. This flow separation only occurs in the FSGe model with $K_{\tau\sigma,o}=0.0$ (top left in Figure 12) due to its nearly two-fold increase in radius. Note, however, that the maximum backflow velocity is only 1.6% of the maximum inlet velocity. This effectively reverses the direction of the local WSS vector $\bm{\tau}_{w}$ acting on the vessel wall. However, in solid G&R models, $\tau_{w}$ can only be estimated as a scalar quantity to inform the WSS stimulus, typically derived from Poiseuille flow (1). It is well established that not only the magnitude but also the pulsatility and direction of WSS are constantly sensed by the endothelium [82, 83]. For example, the directionality of WSS plays an important role in the development of atherosclerosis [84]. A directional metric of WSS has been incorporated in the previous FSG approach of Teixeria et al. [11]. Here, the oscillatory nature of flow is quantified by the wall shear stress aspect ratio, which captures the multi-directionality of flow [27]. In addition, Mohamied et al. [85] proposed a new metric,
@@
to quantify the change of direction of WSS over the course of a cardiac cycle $T$, with vessel normal vector $\bm{n}$ and mean WSS vector $\bm{\tau}_{\text{mean}}$. This metric was shown to correlate better with atherosclerotic lesion prevalence than other metrics like time-averaged or oscillatory WSS index. Our FSGe approach allows directional stimuli for $\Delta\tau_{w}$ by replacing the L2-norm of the WSS vector $\mathbf{\tau}_{w}$ in (17) by a directionality measure like transWSS (41). In the future, incorporating local and directional WSS will enable the testing of hypotheses in FSGe-mediated aneurysm growth and atherosclerotic lesion formation using FSGe. Since current G&R models were designed with the static and scalar nature of reduced fluid dynamics in mind, we expect future biological insights from FSGe to arise from the advancements outlined in this work.